OpenStudy (anonymous):

One number is 2 times another number plus 3. Their sum is 21. Find the numbers. If x represents the smaller number, which of the following equations could be used to solve the problem? 2x + 3 = 21 3x + 3 = 21 3x + 6 = 21

OpenStudy (anonymous):

Let first number be x and the second number be y y = 2x + 3_______1 x + y = 21 y = 21 - x _______2 Substitute 2 into 1 21 - x = 2x + 3 Rearrange the equation ^ 2x + 3 = 21 - x 3x + 3 = 21 THEREFORE, The 2nd option out of the three is correct.